# A Survey on Multilingual Mental Disorders Detection from Social Media Data

### v1

@@ A.1 Methodology details
Initially, 405 papers were retrieved through a database search across the ACL Anthology, ACM Digital Library, IEEE Xplore, Springer Nature Link, ScienceDirect, and Google Scholar. After screening and assessing their eligibility, 108 papers were included in this survey. The PRISMA flow diagram is presented in Figure 3.
Figure 3: PRISMA flow diagram for our review.
@@ A.2 Rankings of the publication venues for the multilingual datasets
Maghraby and Ali (2022) Arabic depression Twitter PHQ-9 PHQ-9 symptoms 1.2K posts FREE TF-IDF, RF F1-score: 98%
AraDepSu Hassib et al. (2022) Arabic depression, suicide Twitter Manual annotation Depression, depression with suicidal ideation, or non-depression 20K posts UNK MARBERT Accuracy: 91.20%, F1-score: 88.75%
@@
Baghdadi et al. (2022) Arabic suicide Twitter Manual annotation Binary 2K posts FREE AraBERT Accuracy: 96.06%, F1-score: 95.86%
Abdulsalam et al. (2024) Arabic suicide Twitter Manual annotation Binary 5.7K posts UNK AraBERT Accuracy: 91%, F1-score: 88%
@@
BanglaSPD Islam et al. (2022) Bengali suicide Facebook Manual annotation Binary 1.7K posts UNK fastText, CNN-BiLSTM Accuracy: 61%, F1-score: 61%
@@
de Carvalho et al. (2020) Brazilian Portuguese suicide Twitter Manual annotation Possibly/ Strongly concerning, Safe to ignore 2.4K posts UNK BERT-Portuguese F1-score: 79%
@@
Oliveira et al. (2024) Brazilian Portuguese suicide Twitter Manual annotation Binary 3.7K posts FREE BERTimbau Accuracy: 96%
Gao et al. (2019) Cantonese suicide Youtube Manual annotation Binary 5K posts UNK Word2vec, LSTM Geometric mean of accuracies: 84.5%
Zhang et al. (2014) Chinese suicide Sina Weibo SPS SPS score 697 users UNK LIWC, LR RMSE: 11
Huang et al. (2015) Chinese suicide Sina Weibo Manual annotation Binary 7.3K posts UNK Topic modeling, LibSVM 80.0%
Cheng et al. (2017) Chinese suicide Sina Weibo Suicide Probability Scale (SPS), DASS-21 Binary 974 users UNK LIWC, SVM AUC: 0.61%
@@
Cao et al. (2019) Chinese suicide Sina Weibo Manual checking of self-report and/or appartenence to a suicide-related community Binary 7K users DUA fastText, RNN Accuracy: 91%
@@
Peng et al. (2019) Chinese depression Sina Weibo Manual annotation Binary 387 users UNK TF-IDF, SVM 83.46%
Huang et al. (2019) Chinese suicide Sina Weibo Manual annotation Binary 18.5K posts UNK LIWC, Dictionary, LR, DT, SVM F1-score: 0.88%
@@
Sun et al. (2022) Chinese suicide, depression Sina Weibo BDI, SDS, Manual annotation Binary / Possibly/Strongly concerning, Safe to ignore 203 users, 1.2K posts UNK Gradient Boosting Accuracy: 82.4%
@@
Wu et al. (2023) Chinese suicide Dcard and PTT Manual annotation Risk levels 2K posts UNK - -
@@
Wang et al. (2024) Chinese depression Sina Weibo Manual annotation Binary 14.8K users AUTH Multimodal features, DNN F1-score: 89.15%
@@
Desmet and Hoste (2014) Dutch suicide Online forums Manual annotation Fine-grained labels 1.3K posts UNK BOW, SVM F1-score: 85.6%
Desmet and Hoste (2018) Dutch suicide Online forums Manual annotation Fine-grained labels 10K posts UNK BOW, Topic modeling, LibSVM F1-score: 92.69%
@@
Astoveza et al. (2018) Filipino, Taglish suicide Twitter Manual annotation Binary 2.1K posts UNK BOW ,MLP Accuracy: 77.9%
@@
SMHD-GER Zanwar et al. (2023) German depression, ADHD, anxiety, bipolar, OCD, PTSD, schizophrenia Reddit Manual annotation Labels for multiple disorders 28K posts DUA LIWC, BiLSTM F1-score: 52.22%
@@
Agarwal and Dhingra (2021) Code-Mixed Hindi-English suicide Reddit Subreddit membership Binary 6.4K posts FREE Indic BERT Accuracy: 98.54%
Oyong et al. (2018) Indonesian depression Twitter Manual annotation Binary 55 users UNK Hand-crafted depression score F1-score: 0.50%
@@
Wang et al. (2023) Japanese suicide Twitter Manual annotation Binary 30K posts N/A
Lee et al. (2020) Korean suicide Naver Cafe Membership in a forum Binary 31K posts UNK Word2Vec, RNN Accuracy: 87.49%
Park et al. (2020) Korean suicide Online forums Manual annotation Risk levels 2.7K posts AUTH XLM-R Accuracy: 88%
Kim et al. (2022a) Korean suicide Twitter Manual annotation Binary 20K posts, 414 users UNK
Kim et al. (2022b) Korean depression Online forums PHQ-9, Manual annotation PHQ-9 score, PHQ-9 symptoms 60 users, 28K posts UNK BERT-based Accuracy: 68.3%
Jung et al. (2023) Korean suicide Twitter Manual annotation Binary 20k posts UNK Metadata, word count, XGBoost F1-score: 83.57%
@@
Uddin (2022) Norwegian depression Online forums Manual annotation Binary 21.8K posts UNK TF-IDF, LSTM Accuracy: 99%
Uddin et al. (2022) Norwegian depression Online forums Manual annotation Binary 30K posts UNK Hand-crafted depression features; LSTM Accuracy: 99%
@@
Rehmani et al. (2024) Roman Urdu depression Facebook Manual annotation Depression severity 3K posts AUTH SVM Accuracy: 84%
Mohmand et al. (2024) Roman Urdu depression Twitter Keywords-based annotations + Expert review Depression severity 25K posts FREE Transfer learning; BERT Accuracy: 99%
Stankevich et al. (2019) Russian depression VKontakte BDI BDI score 531 users UNK Psycholinguistic Markers F1 Score: 66%
@@
Rathnayake and Arachchige (2021) Sinhala depression Twitter, Facebook Manual annotation Binary 1K posts UNK KNN Accuracy: 70%
@@
Herath and Wijayasiriwardhane (2024) Sinhala suicide Facebook Manual annotation Binary 300 posts UNK Naive Bayes Accuracy: 79%
@@
SAD López-Úbeda et al. (2019) Spanish anorexia Twitter Hashtags Binary 5.7K posts FREE SVM Accuracy: 91.6%
Valeriano et al. (2020) Spanish suicide Twitter Manual annotation Binary 2K posts FREE Word2Vec; LR Accuracy: 79%
Ramírez-Cifuentes et al. (2020) Spanish suicide Twitter Manual annotation Binary 252 users N/A
Ramírez-Cifuentes et al. (2021) Spanish anorexia Twitter Manual annotation Anorexia, control, under treatment, recovered, doubtful 645 users N/A
Villa-Pérez et al. (2023) Spanish, English depression, ADHD, anxiety, ASD, bipolar, eating disorders, OCD, PTSD, schizophrenia Twitter Self-disclosure Labels for multiple disorders 6K users DUA N-Grams; XGBoost AUC: 71.2%
MentalRiskES Romero et al. (2024) Spanish depression, anxiety, suicide, eating disorders Telegram Manual annotation Binary + suffer + in favour (sf), suffer + against (sa), suffer + other (so) for Depression 1.2K users AUTH Social media text; mDeBERTa F1 Score: 46%
Cremades et al. (2017) Spanish, English suicide Facebook, Twitter, Blogspot, Reddit, Pinterest Manual annotation Binary 97 posts FREE
Coello-Guilarte et al. (2019) Spanish, English depression Twitter Self-disclosure Binary 316 users FREE BA-LIWC F1 Score: 65%
@@
Mahasiriakalayot et al. (2022) Thai depression Twitter Manual annotation Depression symptoms 3.1K posts UNK LSTM Accuracy: 89.17%
Boonyarat et al. (2024) Thai suicide Twitter Manual annotation Binary + 6 emotions 2.4K posts FREE Linguistic features; BERT F1-score: 90%
Benjachairat et al. (2024) Thai suicide Twitter Manual annotation C-SSRS Labels 5.1K posts AUTH Text features; LSTM F1-score: 93.88%

### v2

@@ A.1 Methodology details
The search terms used in our survey are the following: (“mental disorder” OR “mental illness” ” OR “mental health” OR “depression” OR “anxiety” OR “bipolar” OR “ptsd” OR “post traumatic stress disorder” OR “suicide” OR “suicide ideation” OR “ocd” OR “obsessive compulsive disorder” OR “schizophrenia” OR “eating disorder” OR “anorexia” OR “bulimia”) AND “social media” AND (“multilingual” OR the list of languages provided in Joshi et al. (2020)).
@@ A.2 Rankings of the publication venues for the multilingual datasets
Maghraby and Ali (2022) Arabic depression Twitter PHQ-9 PHQ-9 symptoms 1.2K posts FREE - -%
AraDepSu Hassib et al. (2022) Arabic depression, suicidal ideation Twitter Manual annotation Depression, depression with suicidal ideation, or non-depression 20K posts UNK MARBERT Accuracy: 91.20%, F1-score: 88.75%
@@
Baghdadi et al. (2022) Arabic suicidal ideation Twitter Manual annotation Binary 2K posts FREE AraBERT Accuracy: 96.06%, F1-score: 95.86%
Abdulsalam et al. (2024) Arabic suicidal ideation Twitter Manual annotation Binary 5.7K posts UNK AraBERT Accuracy: 91%, F1-score: 88%
@@
BanglaSPD Islam et al. (2022) Bengali suicidal ideation Facebook Manual annotation Binary 1.7K posts UNK fastText, CNN-BiLSTM Accuracy: 61%, F1-score: 61%
@@
de Carvalho et al. (2020) Brazilian Portuguese suicidal ideation Twitter Manual annotation Possibly/ Strongly concerning, Safe to ignore 2.4K posts UNK BERT-Portuguese F1-score: 79%
@@
Oliveira et al. (2024) Brazilian Portuguese suicidal ideation Twitter Manual annotation Binary 3.7K posts FREE GPT-4 F1-score: 98%
Gao et al. (2019) Cantonese suicidal ideation Youtube Manual annotation Binary 5K posts UNK Word2vec, LSTM Geometric mean of accuracies: 84.5%
Zhang et al. (2014) Chinese suicidal ideation Sina Weibo SPS SPS score 697 users UNK LIWC, LR RMSE: 11
Huang et al. (2015) Chinese suicidal ideation Sina Weibo Manual annotation Binary 7.3K posts UNK Topic modeling, LibSVM F1-score: 80.0%
Cheng et al. (2017) Chinese suicidal ideation Sina Weibo Suicide Probability Scale (SPS), DASS-21 Binary 974 users UNK LIWC, SVM AUC: 61%
@@
Cao et al. (2019) Chinese suicidal ideation Sina Weibo Manual checking of self-report and/or appartenence to a suicide-related community Binary 7K users DUA fastText, RNN F1-score: 90.92%
@@
Peng et al. (2019) Chinese depression Sina Weibo Manual annotation Binary 387 users UNK TF-IDF, SVM F1-score: 76.12%
Huang et al. (2019) Chinese suicidal ideation Sina Weibo Manual annotation Binary 18.5K posts UNK LIWC, Dictionary, LR, DT, SVM F1-score: 88%
@@
Sun et al. (2022) Chinese suicidal ideation, depression Sina Weibo BDI, SDS, Manual annotation Binary / Possibly/Strongly concerning, Safe to ignore 203 users, 1.2K posts UNK Gradient Boosting F1-score: 78.90%
@@
Wu et al. (2023) Chinese suicidal ideation Dcard and PTT Manual annotation Risk levels 2K posts UNK - -
@@
Wang et al. (2024a) Chinese depression Sina Weibo Manual annotation Binary 14.8K users AUTH Multimodal features, DNN F1-score: 89.15%
@@
Desmet and Hoste (2014) Dutch suicidal ideation Online forums Manual annotation Fine-grained labels 1.3K posts UNK BOW, SVM F1-score: 85.6%
Desmet and Hoste (2018) Dutch suicidal ideation Online forums Manual annotation Fine-grained labels 10K posts UNK BOW, Topic modeling, LibSVM F1-score: 92.69%
@@
Astoveza et al. (2018) Filipino, Taglish suicidal ideation Twitter Manual annotation Binary 2.1K posts UNK BOW, MLP Accuracy: 77.9%
@@
SMHD-GER Zanwar et al. (2023) German depression, ADHD, anxiety, bipolar, OCD, PTSD, schizophrenia Reddit Manual annotation Labels for multiple disorders 28K posts DUA LIWC, BiLSTM F1-score: 50.89%
@@
Agarwal and Dhingra (2021) Code-Mixed Hindi-English suicidal ideation Reddit Subreddit membership Binary 6.4K posts FREE Indic BERT Accuracy: 98.54%
Oyong et al. (2018) Indonesian depression Twitter Manual annotation Binary 55 users UNK Hand-crafted depression score F1-score: 50%
@@
Wang et al. (2023) Japanese suicidal ideation Twitter Manual annotation Binary 30K posts N/A
Lee et al. (2020) Korean suicidal ideation Naver Cafe Membership in a forum Binary 31K posts UNK Word2Vec, RNN F1-score: 87.49%
Park et al. (2020) Korean suicidal ideation Online forums Manual annotation Risk levels 2.7K posts AUTH KoBERT F1-score: 88%
Kim et al. (2022a) Korean suicidal ideation Twitter Manual annotation Binary 20K posts, 414 users UNK
Kim et al. (2022b) Korean depression Online forums PHQ-9, Manual annotation PHQ-9 score, PHQ-9 symptoms 60 users, 28K posts UNK BERT-based F1-score: 93%
Jung et al. (2023) Korean suicidal ideation Twitter Manual annotation Binary 20k posts UNK Metadata, word count, XGBoost F1-score: 83.57%
@@
Uddin (2022) Norwegian depression Online forums Manual annotation Binary 21.8K posts UNK TF-IDF, LSTM F1-score: 97%
Uddin et al. (2022) Norwegian depression Online forums Manual annotation Binary 30K posts UNK Hand-crafted depression features; LSTM F1-score: 98%
@@
Rehmani et al. (2024) Roman Urdu depression Facebook Manual annotation Depression severity 3K posts AUTH SVM F1-score: 70.3%
Mohmand et al. (2024) Roman Urdu depression Twitter Keywords-based annotations + Expert review Depression severity 25K posts FREE Transfer learning; BERT F1-score: 99%
Stankevich et al. (2019) Russian depression VKontakte BDI BDI score 531 users UNK Psycholinguistic Markers F1 Score: 65%
@@
Rathnayake and Arachchige (2021) Sinhala depression Twitter, Facebook Manual annotation Binary 1K posts UNK KNN F1-score: 68.5%
@@
Herath and Wijayasiriwardhane (2024) Sinhala suicidal ideation Facebook Manual annotation Binary 300 posts UNK SVM F1-score: 76%
@@
SAD López-Úbeda et al. (2019) Spanish anorexia Twitter Hashtags Binary 5.7K posts FREE SVM F1-score: 91.6%
Valeriano et al. (2020) Spanish suicidal ideation Twitter Manual annotation Binary 2K posts FREE Word2Vec; LR F1-score: 79%
Ramírez-Cifuentes et al. (2020) Spanish suicidal ideation Twitter Manual annotation Binary 252 users N/A Psychological features, LR F1-score: 80%
Ramírez-Cifuentes et al. (2021) Spanish anorexia Twitter Manual annotation Anorexia, control, under treatment, recovered, doubtful 645 users N/A CNN, LR F1-score: 98%
Villa-Pérez et al. (2023) Spanish, English depression, ADHD, anxiety, ASD, bipolar, eating disorders, OCD, PTSD, schizophrenia Twitter Self-disclosure Labels for multiple disorders 6K users DUA N-Grams; XGBoost F1-score: 82.4%
MentalRiskES Romero et al. (2024) Spanish depression, anxiety, suicidal ideation, eating disorders Telegram Manual annotation Binary + suffer + in favour (sf), suffer + against (sa), suffer + other (so) for Depression 1.2K users AUTH Social media text; mDeBERTa F1 Score: 46%
Cremades et al. (2017) Spanish, English suicidal ideation Facebook, Twitter, Blogspot, Reddit, Pinterest Manual annotation Binary 97 posts FREE
Coello-Guilarte et al. (2019) Spanish, English depression Twitter Self-disclosure Binary 316 users FREE BOW, SVM F1 Score: 78%
@@
Mahasiriakalayot et al. (2022) Thai depression Twitter Manual annotation Depression symptoms 3.1K posts UNK LSTM F1-score: 81.06%
Boonyarat et al. (2024) Thai suicidal ideation Twitter Manual annotation Binary 2.4K posts FREE Linguistic features; BERT F1-score: 93%
Benjachairat et al. (2024) Thai suicidal ideation Twitter Manual annotation C-SSRS Labels 5.1K posts AUTH Text features; LSTM F1-score: 93.88%